# 2Huawei Noah’s Ark Lab, London, UK

### v1

@@ 2.3 Affordance-Oriented Robotic Manipulation
Recent work voxposer ; lerf-togo ; wang2023d3fields ; shen2023F3RM ; tsagkas2024click has achieved task-oriented grasping with the help of vision-language models (VLMs) and large language models (LLMs) clip ; dinov2 ; dino ; stable-diff ; gpt4v ; gpt4 . This involves using LLMs to infer parts to be grasped based on task instructions, and VLMs to precisely locate specific parts. Although this pipeline can produce effective and accurate manipulation, the reasoning process requires extra prompt engineering and is time-consuming. In contrast, less attention has been paid to more efficient affordance-oriented grasping that can derive graspable and functional areas without explicitly specifying corresponding object parts. For example, if tasked with slicing bread, an affordance-oriented system can deduce that the serrated edge of a bread knife is appropriate for slicing, while the handle is the correct part to be grasped. One major barrier is the lack of large-scale affordance datasets, compounded by the difficulty in unifying different datasets due to varying annotation standards. Additionally, a recent study ooal highlighted the insufficient granularity of affordance features extracted from existing foundation models, especially when language instructions are involved. For instance, they would associate “cutting” with the entire knife rather than its blade.
To overcome the data limitations, recent methods such as VRB vrb and Robo-ABC robo-abc have explored the potential of learning affordances from human videos for robotic manipulation. However, VRB generates only coarse Gaussian heatmaps and necessitates additional policy learning to deploy the affordance model in real robots. Robo-ABC relies on point correspondences that can be noisy and susceptible to background variations, and do not necessarily lead to reliable grasp poses. In contrast, we propose a more effective and robust strategy to extract precise affordance masks, and introduce an effective model for affordance learning. When combined with a handful of basic pre-recorded motion primitives (as in dipalo2024on ; dipalo2024dinobot ) and grasp pose detection models contact-grasp ; anygrasp ; graspnet , our method enables the utilization of both graspable and functional affordances for robotic manipulation.
3 Method
Abbildung 2: Illustration of the data collection process from egocentric videos. First, graspable points (depicted in purple) are localized from clips of hand-object interaction and then projected to pre-contact frame by homography. Next, functional points (depicted in green) are identified from tool-object interactions and mapped to the pre-contact frame of hand-object interaction through point correspondence. Lastly, these points are used as prompts for the SAM to obtain affordance masks.
In this study, our goal is to develop a holistic system that covers data collection, model learning, and robot deployment. To this end, we first develop an automated pipeline to collect images and related affordance annotations from human videos. Next, we propose an effective affordance learning model termed Geometry-guided Affordance Transformer (GAT). GAT is based on a vision transformer vit using DINOv2 as the encoder dinov2 , and it incorporates a depth feature injector and LoRA layers to facilitate training. Finally, we introduce Aff-Grasp that couples the trained model with an off-the-shelf grasp generation model to achieve affordance-oriented manipulation. Given a task and a cluttered scene, the robot can identify the object with the appropriate affordance, grasp the correct part, and use its functional component on the target object. In Section 3.1, we describe how affordance data are collected from large-scale egocentric videos of human interactions. In Section 3.2, we elaborate on the design of GAT that enables effective affordance learning from collected data. Lastly, in Section 3.3, we explain Aff-Grasp, detailing how it yields affordance-oriented grasp poses for robotic manipulation.
3.1 Data Collection from Egocentric Videos
Given an egocentric video of a human interacting with an object, our aim is to first locate contact points. Human-object interaction videos can generally be categorized into two types: hand-object interaction and tool-object interaction. In hand-object interaction, contact points indicate where the human grasps the object. In tool-object interaction, contact points reveal which part of the tool is used to interact with the target object. These points represent sparse graspable and functional areas of an object, carrying rich affordance information. As shown in Figure 2, we propose a pipeline to automatically collect these points without manual annotation. The collected points then serve as prompts to produce precise segmentation masks using the Segment Anything Model (SAM) sam .
3.1.1 Graspable Point Localization
Egocentric videos like Epic-Kitchens epic and Ego4D ego4d include timestamped narrations that describe actions and their respective start and end times. Based on these narrations, we first retrieve hand-object interaction clips (associated with actions such as “take” or “hold”) and employ a hand-object detector ho-detector to generate contact states and hand-object bounding boxes for all frames. Next, we use labeled timestamps to extract the contact frame, which is typically annotated as the start of an action. We conduct an additional hand segmentation in this frame using the hand box as a prompt via EfficientSAM eff-sam . We then locate the intersection region of the hand mask and object bounding box to sample $n$ contact points $P=\{p_{1},p_{2},...,p_{n}\}$. However, the sampled points are often occluded by hands, and therefore do not accurately represent the graspable affordance area of the object. To collect clean object images free of occlusion, it is necessary to identify the pre-contact frame, i. e., the last frame where the object is fully visible before contact occurs. We utilize the contact states to detect the frame that is closest to the contact frame but without hand-object contact, designating this as the pre-contact frame. Since human motion between adjacent frames is minimal, we follow a similar pipeline to previous studies joint_hand_hotspot ; robo-abc to project the average position of sampled graspable points to the pre-contact frame by computing a homography transformation, as illustrated by the purple dashed line in Figure 2. Specifically, correspondences between these two frames are detected using the SURF descriptor surf , and the homography is then estimated by sampling at least four pairs of points with the RANSAC ransac algorithm to maximize the number of inliers.
3.1.2 Functional Point Localization
To localize functional points, we first retrieve the relationship between objects and affordances from existing affordance datasets. We then extract related tool-object interaction clips from video narrations based on the object-affordance relationship. For example, most datasets associate affordances “cut” and “grasp” with a knife. After localizing the graspable points from a clip showing a person grasping a knife, we then retrieve the next nearest clip depicting a cutting action with the knife.
However, the tool is often heavily occluded or invisible in the contact frame. Similar to graspable point localization, we need to find the pre-contact frame that shows minimal or no intersection between the tool and the target object. To achieve this, we first employ the same combination of the hand-object detector and EfficientSAM to obtain the bounding box and mask of the hand-held tool. Next, we use an open-vocabulary object segmentation model, GroundedSAM groundedsam , to segment the target object. We then measure the Intersection over Union (IoU) between tool and object bounding boxes in frames prior to the contact frame until the IoU is below a preset threshold. Lastly, we calculate the point distances between masks in this pre-contact frame and extract the point within the tool mask that has the shortest distance to all points in the object mask. An erosion operation is applied to the object mask to ensure that the functional point is inside the object.
Nonetheless, not all object categories have related action clips in the narrations. In such cases, we use the farthest sampling to determine the functional points based on the distance to the grasp points. This simple method also produces accurate functional points, as most tools are designed with graspable and functional parts distributed at opposite ends.
Abbildung 3: The architecture of GAT. It consists of a DINOv2 image encoder, a depth feature injector, an embedder, and LoRA layers. The model performs segmentation by computing cosine similarity between upsampled features and learnable or CLIP text embeddings.
3.1.3 Training Data Generation
After extracting functional points, we first project these points to the pre-contact frame of the hand-object interaction clip where we infer the graspable points. Since the object category remains the same, we compute the point correspondence within object bounding boxes using foundation model features deepvit , which map the functional point from the tool-object pre-contact frame to the hand-object pre-contact frame (illustrated by the green dashed line in Figure 2). We then label the graspable points as positive and the functional points as negative to obtain the graspable affordance mask. Conversely, the functional affordance mask is generated by reversing the positive and negative labels. Finally, we crop the object images and store them, along with the generated segmentation masks as annotations.
3.2 Geometry-guided Affordance Transformer
The affordance data are collected from egocentric videos without manual labor, but they have two major issues that hinder effective model training. The first issue is the low resolution of the collected images. The object of interests occupy very small areas of the video frames, often resulting in cropped images smaller than 100 pixels in either length or width. The second one is the limited diversity of the training data, characterized by monotonous backgrounds and mostly restricted to indoor scenes. As a result, training a model using typical ImageNet imagenet pre-trained representation leads to fairly poor performance (detailed in Section 4.1.3).
To cope with the above issues, we propose an affordance learning architecture called Geometry-guided Affordance Transformer (GAT). The illustration of GAT is presented in Figure 3. It includes a novel Depth Feature Injector (DFI) that integrates geometric priors into image features using pseudo depth maps, a DINOv2 image encoder as feature extractor, and additional LoRA layers for effective fine-tuning.
3.2.1 Depth Feature Injector
We argue that depth maps introduce rich geometric information that can help with foreground-background and part separation. Additionally, they exclude color information, allowing the model to fully focus on the shape information, which is highly relevant to affordances. For instance, graspable parts often consist of shapes like cylinders or spheres, while parts designed for cutting typically feature sharp edges and flat surfaces.
Specifically, we first obtain pseudo depth maps for each training image with a state-of-the-art depth estimation model Depth-Anything depth-any . During model training, the pseudo depth map is first encoded into feature maps using a stem block, which contains several standard $3\times 3$ convolution layers as in the ResNet resnet . These feature maps are then processed by a $1\times 1$ convolution that transforms the channel dimension to match that of the RGB image features.
$Q=\phi_{q}(F_{i}),\leavevmode\nobreak\ K=\phi_{k}(F_{d}),\leavevmode\nobreak\ V=\phi_{v}(F_{d}),$ (1)
$\hat{F}_{i}=\beta\cdot\text{softmax}(QK^{T}/\sqrt{d_{k}})\cdot V+F_{i},$ (2)
where $\phi$ is a linear transformation, and $d_{k}$ is the dimension of the key acting as a scaling factor. Following vit-adapter , we set a learnable vector $\beta\in\mathbb{R}^{C}$, initialized to 0, to balance the output from the cross-attention layer and the image feature. This strategy prevents the image feature from being excessively affected by the depth feature, making the training process more stable.
We observe that DFI constantly brings improvement, even when integrated solely during training (see Section 4.3). This indicates that it can act as a regularization mechanism during training, and can be discarded during inference to speed up the process.
3.2.2 DINOv2 with Low-Rank Adaptation
We notice that directly training from the typical ImageNet pre-trained representation often leads to inferior results. This can be attributed to two primary reasons: First, affordance segmentation focuses on fine-grained object parts, whereas representations trained for image classification emphasize more on global object features deepvit . Second, ImageNet pre-trained models exhibit limited diversity, making it challenging to handle data from diverse domains. To address this issue, we employ the self-supervised visual foundation model DINOv2, which has been demonstrated to be highly effective for data-limited affordance learning due to its properties of part-aware representation and part-level correspondence ooal . Furthermore, we introduce LoRA lora to fine-tune the model without modifying the parameters of the original DINOv2. This strategy helps adaptation across different domains and prevents overfitting. LoRA was originally developed to fine-tune large language models for different downstream tasks. Specifically, it injects trainable rank decomposition matrices to a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ by $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During training, only $A$ and $B$ are trainable, while $W_{0}$ remains frozen. This incurs minimal computational cost and memory usage. With LoRA layers, the original forward pass $h=W_{0}x$ is modified as:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx.$ (3)
We apply LoRA to all query, key, and value projection layers, and find that this fine-tuning strategy leads to better transfer learning results.
Abbildung 4: The framework of Aff-Grasp. It first employs an open-vocabulary detector to locate all objects within the scene, which are then sent to GAT to dertermine if they possess corresponding affordance required for the task. Afterwards, a 6 DoF grasp generation model, utilizing both the object’s graspable affordance and the depth map, estimates the potential grasp poses. Finally, the robot executes affordance-specific sequential motion primitives to apply the functional part to the target.
3.2.3 Classifier and Loss Functions
To speed up inference time for real-world applications, we avoid adding any complex decoder structures. Instead, we process the output feature with an embedder (an MLP), reshape it, and upsample it by a factor of four to increase the resolution $F_{out}\in\mathbb{R}^{C\times\frac{4H}{p}\times\frac{4W}{p}}$, where $p$ is the patch size. Next, we intialize $M\in\mathbb{R}^{{L\times C}}$ learnable embeddings, where $L$ is the number of affordance categories. We compute the cosine similarity between $M$ and $F_{out}$ to yield the segmentation output, which is then restored to the same size of the input image via bilinear interpolation. Due to the domain gap, we do not add a learnable embedding for the background classification to prevent overfitting. Alternatively, we determine a pixel as background if all its affordance predictions are below a preset threshold $\tau$. Compared to a linear layer and explicit background classifier, the cosine similarity-based segmentation and implicit background prediction are more robust and can effectively improve the performance, as detailed in Section 4.3. In addition, to achieve open vocabulary affordance segmentation, the $M$ can also be replaced with corresponding CLIP text embeddings, as verified in ooal .
Since the collected data are highly unbalanced, we utilize a combination of focal loss focal and dice loss dice as training objectives:
$\begin{split}\mathcal{L}_{focal}=-\frac{1}{n}\sum_{i=1}^{n}\big{[}&(1-\hat{y}_{i})^{\gamma}\cdot\hat{y}_{i}\log(y_{i})\\ &+\hat{y}_{i}^{\gamma}\cdot(1-\hat{y}_{i})\log(1-y_{i})\big{]},\end{split}$ (4)
$\mathcal{L}_{dice}=1-\frac{2\sum^{n}_{i}y_{i}\hat{y}_{i}+\epsilon}{\sum^{n}_{i}y_{i}+\sum^{n}_{i}\hat{y}_{i}+\epsilon},$ (5)
$\mathcal{L}=\alpha\cdot\mathcal{L}_{focal}+\mathcal{L}_{dice},$ (6)
where $n$ is the number of valid pixels in output, $\gamma=2$ is a focusing parameter to balance easy and hard samples, $\epsilon=1$ is a smoothing factor that prevents division by zero and stabilizes the training, and $\alpha$ is a weighting factor to balance loss values.
3.3 Affordance-Oriented Robotic Manipulation
@@
Our ultimate goal is affordance-oriented robotic manipulation, where given a task and a cluttered scene, the robot can select the object that possesses the related affordance, grasp the correct part, and apply the functional part to the target object to perform desired actions. To achieve this, we propose Aff-Grasp, which integrates GAT to achieve affordance segmentation and transforms the visual affordance to available grasp poses. The framework of Aff-Grasp is shown in Figure 4. Given a task consisting of a verb and a target, such as “cut cake”, it first uses an open-vocabulary object detection model groundingdino to detect the target (cake) and other visible objects. For object other than the target, the input vocabulary is simply set to “objects” for class-agnostic detection. These detected objects are then cropped and sent to GAT to predict affordance areas. The object with the most certain and largest affordance area for the required action (cut) is identified. After that, we extract the graspable affordance area of this object, and generate potential grasp poses within the area, i. e., the knife handle. For grasp pose generation, we select Contact-GraspNet contact-grasp that can produce dense grasp proposals within a specified mask area. Once the object is grasped and lifted, we execute affordance-specific sequential motion primitives to apply the functional affordance area to the target object to complete the task. For the handover task, Aff-Grasp is instructed to find available grasp proposals within the functional affordance area and then pass the graspable part to the human hand.
When CLIP text embeddings are used as classifiers, the action required for a task can be transformed into text embeddings for open-vocabulary affordance segmentation ov-aff-3d ; ooal . Therefore, unseen affordance vocabularies can also be used at inference, enhancing the model’s adaptability and versatility. Overall, Aff-Grasp seamlessly integrates affordance prediction and grasp pose generation to enable robust and flexible robotic manipulation in diverse and cluttered environments.
@@ 4 Experiments
In this section, we present experiments from both vision and robot perspectives. Section 4.1 details the vision experiments, where we propose a dataset for evaluation and compare GAT with state-of-the-art models. Section 4.2 describes robotic experiments and compares our grasping framework Aff-Grasp with two related methods that acquire affordance knowledge from human-object interactions. Section 4.3 presents ablation studies that examine the design choices of GAT.
Tabelle 1: Statistics on the number of images for each object on the AED.
Total knife cup scissors hammer fork screwdriver spatula ladle pan shovel spoon drill trowel
721 156 95 78 78 72 59 46 46 31 22 18 10 10
@@ 4.1 Vision Experiments
4.1.1 Evaluation Dataset
To evaluate the effectiveness of GAT, we require a diverse and challenging affordance dataset that has consistent object and affordance categories with the collected training data. After carefully inspecting existing datasets, we found that most of them are not compatible with our evaluation requirements. Many datasets either have a small number of categories aff_with_CNN_iit ; learn2act_properly or are collected in the lab environment with limited diversity tool_parts_umd ; cad120 . Some datasets contain a large number of images but have coarse keypoint-based annotations ag_from_exocentric_imgs ; demo2vec or small resolutions aff_low_resolution_dataset . Therefore, we create an Affordance Evaluation Dataset (AED) by manually annotating 721 images collected from several existing affordance datasets and internet resources. AED contains 13 object categories and 8 affordance classes. In Figure 5, we present examples from existing affordance datasets alongside AED, highlighting the necessity for a new evaluation dataset. The statistics regarding the number of images per object category are listed in Table 1.
Abbildung 5: Examples from existing affordance datasets. UMD tool_parts_umd is collected in fixed lab environment with coarse annotations. RGBD-AFF aff_low_resolution_dataset has very low resolution and clean background. IIT-AFF aff_with_CNN_iit includes humans and occluded parts are also annotated. AGD20K ag_from_exocentric_imgs is annotated with keypoints and transformed to coarse heatmaps with a Gaussian kernel. In contrast, AED contains natural images with pixel-wise annotations.
4.1.2 Implementation Details
All experiments are conducted on two GeForce RTX 3090 GPUs using the Adamw adamw optimizer, with a learning rate of $1\mathrm{e}{-3}$ and batch size 8 for 15 epochs. DINOv2-base is used as the feature extractor. For data collection from egocentric videos, we utilize narratives to filter irrelevant objects and only collect the 13 objects contained in the evaluation dataset. Collected images are first resized to $476\times 476$ and then randomly cropped to $448\times 448$. Both horizontal and vertical flipping are used for data augmentation. Three metrics, mean intersection-over-union (mIoU), F1-score (F1), and accuracy (Acc), are adopted for evaluation.
Abbildung 6: Qualitative comparison between our approach and other segmentation models on the AED.
4.1.3 Quantitative and Qualitative Comparisons
Table 2 shows the results of different state-of-the-art segmentation approaches on the proposed AED. They use either pre-trained ImageNet backbones to extract feature maps, or obtains representations from visual foundation models like CLIP clip and DINOv2 dinov2 . Thus, we divide these models into two sections based on the pre-training strategies.
Tabelle 2: Quantitative comparison on the AED.
@@ 4.1.3 Quantitative and Qualitative Comparisons
ImageNet DeepLabV3+ deeplabv3+ 13.46 22.27 23.05
PSPNet pspnet 16.90 27.32 26.46
SegFormer segformer 23.72 36.86 37.19
Foundation Models ZegCLIP zegclip 18.33 26.41 25.55
DINOv2 dinov2 46.16 62.49 63.61
ViT-Adapter vit-adapter 48.36 64.66 65.80
OOAL ooal 52.72 68.70 65.79
@@
For ImageNet pre-trained models, we employ classical CNN segmentation models such as DeepLabV3+ deeplabv3+ and PSPNet pspnet , as well as a transformer-based segmentation model SegFormer segformer . For visual foundation-based models, we choose ZegCLIP zegclip , DINOv2 dinov2 , ViT-Adapter vit-adapter , and OOAL ooal to compare with GAT, as they represent the state-of-the-art in leveraging visual foundation models for semantic or affordance segmentation tasks. We notice that methods using pre-trained ImageNet backbones generally produce much inferior results compared to those based on foundation models, confirming the huge domain gap between training and evaluation sources. Among the foundation model based approaches, our model GAT significantly outperforms the second best counterpart, OOAL, achieving higher performance in mIoU, F1-score, and accuracy by 15.9%, 12.39%, and 17.72%, respectively.
Figure 6 depicts the qualitative comparison between our methods and other models. We find that models like PSPNet and Segformer often yield incomplete or incorrect affordance predictions, which may result from the low diversity in the pre-trained ImageNet representation. On the other hand, most models based on DINOv2 can coarsely generate correct affordance prediction map, but often suffer from incomplete part activation and noisy segmentation around object boundaries. In contrast, the results from GAT are part-focused, exhibit well-preserved boundary segmentation, and are capable of handling complex objects like drills. Notably, our results stand out from other counterparts when dealing with images containing multiple objects, as shown in the 1st, 3rd, 6th, and 8th examples in Figure 6.
In addition, we perform a further qualitative comparison on unseen objects to explore the models’ generalization ability. As displayed in Figure 7, novel objects such as shears, saw, bowl, and sword are used. It is apparent that Segformer cannot make accurate affordance predictions for these objects. OOAL demonstrates acceptable potential on unseen objects but often produce less confident and inconsistent results. In comparison, GAT shows excellent performance on these out-of-distribution objects with much more complete segmentation maps.
Abbildung 7: Qualitative comparison on unseen objects.

### v2

@@ 3.3 Affordance-Oriented Robotic Manipulation
Our ultimate goal is affordance-oriented robotic manipulation, where given a task and a cluttered scene, the robot can select the object that possesses the related affordance, grasp the correct part, and apply the functional part to the target object to perform desired actions. To achieve this, we propose Aff-Grasp, which integrates GAT to achieve affordance segmentation and transforms the visual affordance to available grasp poses. The framework of Aff-Grasp is shown in Fig. 4. Given a task consisting of a verb and a target, such as “cut cake”, it first uses an open-vocabulary object detection model [29] to detect the target (cake) and other visible objects. For objects other than the target, the input vocabulary is simply set to “objects” for class-agnostic detection. These detected objects are then cropped and sent to GAT to predict affordance areas. The object with the most certain affordance area for the required action (cut) is identified. After that, we extract the graspable affordance area of this object, i.e., the knife handle, and generate potential grasp poses within it. For grasp pose estimation, we select Contact-GraspNet [49] that can produce dense grasp proposals within the specified mask area. The one with the highest score is then chosen as the execution pose. Once the object is grasped and lifted, we execute affordance-specific sequential motion primitives to apply the functional affordance area to the target object to complete the task. For the handover task, Aff-Grasp is instructed to find available grasp proposals within the functional affordance area and then pass the graspable part to the human hand. When CLIP text embeddings are used as classifiers, the action required for a task can be transformed into text embeddings for open-vocabulary affordance segmentation [38, 24]. Therefore, unseen affordance vocabularies can also be used at inference, enhancing the model’s adaptability and versatility.
@@ 4 Experiments
In this section, we present experiments from both vision and robot perspectives, along with ablation studies that examine the design choices of GAT. Note that all evaluations are conducted in a zero-shot or cross-domain setting, since the test data and robot setup contain novel objects or affordances, and differ greatly from the training data collected from egocentric videos. As a result, only models with strong generalization ability can achieve high performance. Implementation details are provided in the supplementary.
@@ 4.1 Vision Experiments
Evaluation Dataset. To evaluate the effectiveness of GAT, a diverse and challenging affordance dataset that has consistent object and affordance categories with the collected training data is needed. After carefully inspecting existing datasets, we found that most of them are not compatible with our evaluation requirements. Many datasets either have a small number of categories [37, 6] or are collected in the lab environment with limited diversity [35, 22]. Some datasets contain a large number of images but have coarse keypoint-based annotations [31, 14] or small resolutions [20]. Therefore, we create the Affordance Evaluation Dataset (AED) by manually annotating 721 images collected from several existing affordance datasets and internet resources. It contains 13 object categories and 8 affordance classes. See supplementary for more details on AED.
@@
ImageNet
DeepLabV3+ [4] 13.46 22.27 23.05
PSPNet [61] 16.90 27.32 26.46
SegFormer [54] 23.72 36.86 37.19
Foundation
Models
ZegCLIP [62] 18.33 26.41 25.55
DINOv2 [42] 46.16 62.49 63.61
ViT-Adapter [5] 50.86 66.88 65.21
OOAL [24] 54.82 70.58 68.00
@@
Table 1: Quantitative comparison on the AED.
Quantitative and Qualitative Comparisons. Table 1 shows the results of different segmentation approaches on the proposed AED. They use either pre-trained ImageNet backbones to extract feature maps, or representations from visual foundation models like CLIP [43] and DINOv2 [42]. Thus, we categorize these models into two groups based on their pre-training strategies. For ImageNet pre-trained models, we employ classical CNN segmentation models such as DeepLabV3+ and PSPNet, and a transformer-based model SegFormer. For visual foundation-based models, we choose ZegCLIP, DINOv2, ViT-Adapter, and OOAL to compare with GAT, as they represent the state-of-the-art in leveraging visual foundation models for semantic or affordance segmentation tasks. We notice that methods using pre-trained ImageNet backbones generally produce much inferior results compared to those based on foundation models. This confirms the substantial domain gap between training and evaluation sources, and demonstrates that foundation models are more resistant to noisy training samples and have better cross-domain capabilities.
Figure 5: Qualitative comparison between our approach and other segmentation models on the AED.
Figure 5 depicts the qualitative comparison between our methods and other models. We find that models like PSPNet and Segformer often yield incomplete or incorrect affordance predictions, which may result from the low diversity in the pre-trained ImageNet representation. On the other hand, OOAL can coarsely generate correct affordance prediction map, but often suffer from incomplete part activation and noisy segmentation around object boundaries. In contrast, the results from GAT are part-focused, and exhibit well-preserved boundaries. Notably, our results stand out from other counterparts when dealing with images containing multiple objects. Additional qualitative comparisons, as well as generalization tests on unseen objects and affordances are provided in the supplementary.